OpenStudy (anonymous):

What is the momentum of a 50-kilogram ice skater gliding across the ice at a speed of 5 m/s? (1 point) 10 kg • m/s 500 kg • m/s 50 kg 250 kg • m/s

OpenStudy (anonymous):

momentum = mass*velocity can u proceed from here?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

momentum = 50*5 = 250 kgm/s

OpenStudy (anonymous):

HAPPY BIRTHDAY :)

OpenStudy (anonymous):

Okay old question! But I'll explain it for further reference for other people that need help. :) Multiplying an object's mass (in kilograms) and it's velocity (in meters per second) will give you the momentum. So in order to solve this problem you would take 50 times 5, which equals 250. Therefor your momentum is 250kg ⋅ m/s. kg = kilogram. m/s = meters per second. Repeat this for other problems. :)